OpenStudy (anonymous):

Solve: x^2 − 2x − 24 = 0. x = 4, x = 6 x = 4, x = −6 x = −4, x = 6 x = −4, x = −6

Parth (parthkohli):

Do you know these methods? if any: Completing the square. Quadratic formula factoring.

OpenStudy (anonymous):

\[\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array} \] Use the quadratic formula

OpenStudy (anonymous):

So I would use ax^2 + bx + c

OpenStudy (anonymous):

It's always a sure way to find the answer if you don't know how to do it any other method.

OpenStudy (anonymous):

yes for example a=1

Parth (parthkohli):

Can you tell me what are a, b and c? Let's do it with quadratic formula.

OpenStudy (anonymous):

a = 1, b = 2 and c = 24

OpenStudy (anonymous):

you forgot the negative signs.

OpenStudy (campbell_st):

find the factors of -24 that add to -2.... the hint is the larger factor is negative

Parth (parthkohli):

Hmm. It's actually a = 1, b = -2, c = -24

OpenStudy (anonymous):

not if it's -b * -2

OpenStudy (anonymous):

It converts to a positive

Parth (parthkohli):

How do we have a positive here?

OpenStudy (anonymous):

Nevermind. I'm thinking of something else.

OpenStudy (anonymous):

a = 1 b = -2 and c = -24.

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

Now what?

Parth (parthkohli):

Now plug that into the formula.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

OH

OpenStudy (campbell_st):

factors of - 24 -24 + 1 = -23 -12 + 2 = -10 - 8 + 3 = -5 -6 + 4 = -2 woohooo its -6 and 4 then 0 = (x - 6)(x+4) so the is x = -4, 6

OpenStudy (anonymous):

1x^2 - 2x - 24

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {-(-2) \pm\sqrt{(-2)^2 + 4(1)(-24)} \over 2(1)} }\)

OpenStudy (anonymous):

I don't get it.

Parth (parthkohli):

Can you do it now?

OpenStudy (campbell_st):

why those numbers ... well its because use substitutions 0 = (6 - 6)(6 + 4) or 0 = (-4-6)(-4 +4) = (0) (10) = (-10)(0) = 0 = 0 wow it works

Parth (parthkohli):

\( \color{Black}{\Rightarrow\Large {2 \pm \sqrt{4 - 96} \over 2}}\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow x = \Large{2 \pm \sqrt{-92} \over 2} }\)

OpenStudy (anonymous):

When you use those things it confuses me.

Parth (parthkohli):

Well anyway my first choice would be factoring.

Parth (parthkohli):

If you're not able to factor, then I reckon you the formula.

OpenStudy (anonymous):

Factor each number?

Parth (parthkohli):

Formula confuses everyone :P

Parth (parthkohli):

See what @campbell_st said.

OpenStudy (eyust707):

it should be b^2 - 4ac not b^2 + 4ac

OpenStudy (campbell_st):

PARTH... you have a hugh error i n the discriminant.... it should be 4 + 96 4 x 1 x -24 = 96

OpenStudy (eyust707):

both ways work

Parth (parthkohli):

Yeah, that's what I was thinking

OpenStudy (anonymous):

I see. Could I do it? No. But do I get this problem, sure. I don't know. It takes me 30 min. to an hour for each problem. It sucks.. truly sucks.

OpenStudy (eyust707):

Do you guys mind if I help him out for a minute?

OpenStudy (campbell_st):

careless... work on factors... if its not obvious go to the general quadratic formula

OpenStudy (anonymous):

What do you mean factors

OpenStudy (eyust707):

Whats the next one?

Parth (parthkohli):

\( \color{Black}{\Rightarrow x = \Large{-(-2)\pm\sqrt{100} \over 2}}\) \( \color{Black}{\Rightarrow x = \Large{2 \pm 10 \over 2} }\) \( \color{Black}{\Rightarrow x = 6 \textbf{ or} -4 }\)

OpenStudy (anonymous):

Solve: 0 = 6x2 − 10x − 4.

Parth (parthkohli):

This time do factorisation :P

OpenStudy (anonymous):

Solve: 0 = 6x^2 − 10x − 4.

OpenStudy (anonymous):

No, @ Parth

Parth (parthkohli):

\( \color{Black}{\Rightarrow 6x^2 - 12x + 2x - 4 = 0 }\) Sorry....uh

Parth (parthkohli):

What you do is first you express \(bx\) as a sum of two numbers \(\alpha x\) and \(\beta x\) such that \(\alpha \times \beta = ac\)

Parth (parthkohli):

And, then, you group the first two and last two terms.

OpenStudy (anonymous):

... I don't get it man.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 6x(x - 2) + 2(x - 2) = 0 }\) \( \color{Black}{\Rightarrow (6x + 2)(x - 2) = 0 }\) Using the zero product theorem, either 6x + 2 = 0, or x - 2 = 0

Parth (parthkohli):

Or wait, you may factor further. \( \color{Black}{\Rightarrow 2(3x + 1)(x - 2) = 0 }\)

OpenStudy (campbell_st):

here is a solution to it take 2 out as a common factor 2(3x^2 - 5x - 2) = 0 multiply 3 by -2 a x c 3 x -2 = -6 factors of -6 that give b are -6 + 1 then its  take out any common factors